OpenStudy (anonymous):

Solve for x. 2x^2 – 5x = 12

OpenStudy (anonymous):

you really don't like maths do you..haha

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nope :P I am really bad at it

OpenStudy (anonymous):

start with \[2x^2-5x-12=0\] then factor

OpenStudy (anonymous):

well bring the 12 to the other side so 2x^2-5x-12=0 then use the quadratic formula

OpenStudy (anonymous):

x=4,x=-1.5

OpenStudy (anonymous):

you can use the quadratic formula, but a miracle occurs here and you can factor as \[(x-4)(2x+3)=0\]

OpenStudy (anonymous):

whats the formula for the quadratic formula

OpenStudy (anonymous):